Please indicate all steps taken to solve this question clearly so it is easy to understand. Thank you.

Problem #8: An initial deposit of $7500 1s made nto an account now, and a second deposit of $18000 1s made înto the same account after 15 years. During the first 10 years, the account earns interest quoted at a nominal quarterly discount rate of d%. After 10 years have passed, any money in the account earns a nominal semi-annual interest rate of 6.4%. At the end of 30 years, the total value of the account will be $1 16 173, what is the value of d (to the nearest 0.1%)? d as a percentage, correct to 1 decimal Problem #8: Just Save Your work has been saved! (Back to Admin Page) Submit Problem #8 for Grading Problem #8| Attempt #1 Your Answer: Attempt #2 Attempt #3 | Attempt #4 Attempt #5 Your Mark:
Answer #1
A B C D E F G H I
2
3 The future value in the account can be calculated as follows:
4
5 Future value for $7500 deposited at Year0 =7500*(1+d/4)(10*4)*(1+6.4%/2)(20*2)
6 =26439.25*(1+d/4)40
7
8 Future value for $18,000 deposited at Year 15 =18,000*(1+6.4%/2)(15*2)
9 $46,308.79 =18000*((1+6.4%/2)^(15*2))
10
11 Total Future Value at year 30 =$46,308.79 + 26439.25*(1+d/4)40
12
13 Given the future value $116,173
14
15 Equation can be written as
16 $116,173=$46,308.79 + 26439.25*(1+d/4)40
17
18 Solving the above equation
19 d 9.8% =4*((((D13-D9)/26439.25)^(1/40))-1)
20
21 Hence Value of d is 9.8%
22